OpenStudy (anonymous):

In a new suburb, it is estimated that the number of registered voters will be modeled by the function N(t) = 11 + 9 t^2 - t^3, where t is time in years and N is in thousands. N(4) = ? ________ which represents ? _______(how many voters)

OpenStudy (anonymous):

I wish I could help, man my teacher was just going over that

OpenStudy (anonymous):

N(4) represents the number of thousands of voters in four (4) years N(t) = 11 + 9 t^2 - t^3 N(4) = 11 + 9 *4^2 - 4^3 N(4) = 11 + 9*16 - 64 N(4) = 11 + 144 - 64 N(4) = 155 - 64 N(4) = 91 thousand voters in four (4) years

OpenStudy (anonymous):

N(4) represents the number of thousands of voters in four (4) years \[N(t) = 11 + 9 t^2 - t^3 \]\[N(4) = 11 + 9 *4^2 - 4^3\]\[N(4) = 11 + 9*16 - 64\]\[N(4) = 11 + 144 - 64\]\[N(4) = 155 - 64\]N(4) = 91 thousand voters in four (4) years